OpenStudy (anonymous):

Determine whether the sequence is arithmetic. If it is, find the common difference.

OpenStudy (anonymous):

\[a) \frac{ 1 }{ 2 },\frac{ 1 }{ 4},\frac{ 1 }{ 6 },\frac{ 1 }{ 8 }......\] \[b) 10,20,30,40,......\]

OpenStudy (anonymous):

a is not arithmetic. notice the difference between 1/2 and 1/4 is - 1/4, but the difference between 1/4 and 1/6 is -1/12 b is arithmetic the common difference between each term is 10

OpenStudy (anonymous):

Awesome! Thanks for your help!

OpenStudy (anonymous):

a,b,c,....are in A.P. if 2b=a+c